# Iterative Trajectory Exploration for Multimodal Agents

### v1

@@ 3.4 Preference Tuning
Objective. In one iteration, we may generate multiple tasks and construct preference data for them. After that, we denote the obtained preference data set as $\mathcal{D}=\{(x_{i},s_{i}^{pre},s_{i}^{dis})\}$, and the number of data is $|\mathcal{D}|=d$. We choose the flexible DPO algorithm,
[…]
@@
Efficiency. One issue is the potential low efficiency in such an online self-exploration framework. A key hyperparameter is $d$, that is, how much preference data is collected in one iteration. If we set a large $d$, it means that we collect a large number of preference data for agent tuning in one iteration. This may cause the problem of data bias, as the data required by the controller may change during training, while our data is generated using a fixed controller in the beginning. If we set a small $d$, it means that we collect a small number of preference data for agent tuning in one iteration. This may cause low efficiency, since we need to load and release the model weights (controller, parameter gradients, tools, and verifier) frequently. To balance it, we set $d$ as 500 in practice via empirical exploration.
Training Scheme. The online self-exploration is performed after an SFT stage for controllers, since the effectinveness of the self-exploration framework requires the controller to have the ability to generate accurate solutions. The SFT stage is the same as MAT (Gao et al., 2024b), where 20K trajectories are used to align the agent controller (Qwen2VL-7B in practice) with desirable outputs. In the self-exploration stage, we use preference tuning to update Qwen2-VL using 16K step-wise preference data in all, where $d$ is 500 and there are 4 iterations in total. The preference tuning process is summarized in Algorithm 1.
Algorithm 1 Training process in SPORT.
0:  Seeds of multimodal tasks, initial agent controller $\pi_{\theta}$, and $\pi_{ref}=\pi_{\theta}$. Preference data $\mathcal{D}=\emptyset$.
0:  Updated agent controller $\pi_{\theta^{*}}$
@@
6:           Sample $n$ solutions $\{t_{i}^{1},c_{i}^{1},\cdots,t_{i}^{n},c_{i}^{n}\}$ based on the history $h_{i}$, and execute them to obtain results $\{o_{i}^{1},\cdots,o_{i}^{n}\}$.
7:           Select the best solution $\{t_{i}^{\star},c_{i}^{\star}$}.
8:           Construct $n-1$ preference pairs, and add them into $\mathcal{D}$.
@@ 4.1 Setting
The performance of the proposed SPORT approach is assessed on the GTA (Wang et al., 2024c) and GAIA (Mialon et al., 2023) benchmarks. Results are compared against agents powered by both closed-source models (e.g., GPT-4, GPT-4o, Claude3) and open-source models, including LLaMA-3-70B-instruct (Dubey et al., 2024), Qwen1.5-72B-chat (Bai et al., 2023), LLaVA-NeXT-8B (Liu et al., 2024a), InternVL2-8B (Chen et al., 2024c), Qwen2-VL-7B (Wang et al., 2024d), and MiniCPM-V-8.5B (Yao et al., 2024). Specifically, we perform direct comparisons with leading agents, such as Lego Agent (AgentLego Contributors, 2023), Sibyl Agent (Wang et al., 2024e), and Warm-up Act Agent (Mialon et al., 2023). As a baseline, we use the Huggingface Agent (HF Agent) (HuggingFace Contributors, 2024), which operates with the same toolset as the Sport Agent . We first evaluate these agents on two benchmarks, then assess the quality of produced preference data, and finally show several visualization examples to demonstrate the effectiveness of our method.
We employ the Qwen-2-VL model as the controller. In the training process of our VLM controller, we freeze the vision encoder and visual token compressor, and fine-tune the language model using LoRA (Hu et al., 2022). We set the rank as $32$ and apply LoRA on query, key, and value projection matrices in all self-attention layers. We use the AdamW optimizer with a cosine annealing scheduler. The learning rate is $61.0e-$ 6 and the batch size is $2$ per device. We set the max context window as $10240$ to support complex trajectories of our agent.
@@
Metric. Following existing methods (Wang et al., 2024c; Gao et al., 2024b), we assess agent performance using three key metrics: AnsAcc, ToolAcc, and CodeExec for the GTA benchmark. AnsAcc gauges the accuracy of predicted answers. ToolAcc evaluates the correctness of tool usage and the quality of answer summaries. CodeExec measures the percentage of generated code that executes without errors. In the GAIA benchmark, we focus on measuring AnsAcc at its three levels.
@@ 4.2 GTA Results
The results on the GTA benchmark are shown in Table 1, where key metrics including AnsAcc, ToolAcc, and CodeExec are reported. Our agent surpasses the Lego agent that utilizes closed-source models (e.g., GPT-4 and GPT-4o), as well as the HF agent that uses closed-source models and open-source models (e.g., InternVL2-8B), showcasing the ability of our SPORT Agent to tackle complex tasks with greater efficiency. A comparison between agents through SFT (i.e., T3-Agent) and our SPORT Agent demonstrates the effectiveness of our online self-exploration framework and the advantages of our Step-wise optimization approach. Our method has about $7\%$ improvements on the final accuracy, since it calls more suitable tools ($8\%$ improvements) and reduces code error ($7\%$ improvements). Compared with the HF agent using GPT-4o and GPT-4o mini, our agent achieves higher ToolAcc and comparable CodeExec. This indicates that the proposed SPORT method improves the planning and reasoning capabilities of agents again.
Table 1: Results on the GTA benchmark
Method Controller AnsAcc ToolAcc CodeExec
Lego Agent GPT-4 46.59 - -
Lego Agent GPT-4o 41.52 - -
Lego Agent GPT-3.5-turbo 23.62 - -
Lego Agent Claude3-opus 23.44 - -
Lego Agent Qwen1.5-72B-chat 13.32 - -
Lego Agent LLaMA3-70B-instruct 8.32 - -
HF Agent GPT-4o 57.05 63.41 95.12
HF Agent GPT-4o mini 57.69 56.10 100.00
HF Agent LLaVA-NeXT-8B 14.10 14.97 25.08
HF Agent InternVL2-8B 32.05 36.75 52.18
HF Agent MiniCPM-V-8.5B 33.97 36.59 56.10
HF Agent Qwen2-VL-7B 42.31 44.85 65.19
T3-Agent MAT-MiniCPM-V-8.5B 52.56 65.85 80.49
T3-Agent MAT-Qwen2-VL-7B 53.85 64.63 84.32
Sport Agent Tuned Qwen2-VL-7B 60.26 72.41 91.87
4.3 Ablation
4.3.1 Effectiveness of iteration step size $d$.
We conduct an ablation study on the GTA benchmark to investigate the impact of the iteration step size $d$. We set $d\in\{200,500,1000\}$, adjusting the number of iterations to (5, 2, 1), respectively, to ensure a total of 1000 tasks are processed in each setting. As shown in Table 2, setting $d=500$ yields the best overall performance across all metrics. When $d=1000$, the model sees all tasks in a single pass, which limits adaptability and leads to a slight drop in answer accuracy and execution success. In contrast, using a smaller step size ($d=200$) results in less diverse updates per iteration, which may reduce robustness. These results suggest that a moderate value of $d$ offers a good balance between update frequency and data diversity, leading to more stable and effective training.
Table 2: Ablation on iteration step $d$ in the GTA benchmark
@@ 4.3 Ablation
4.3.2 Applying DPO to MAT Preference Data
To quantify the impact of directly optimizing MAT trajectories with preference-based learning, we perform an ablation by applying DPO to the original MAT-SFT model. Specifically, we treat the MM-Traj data (GPT-4o generated) as “preferred” samples and synthetically generate an equal number of “dispreferred” examples via the MAT-SFT model, matching the total volume of our SPORT preference data. We then fine-tune MAT-SFT using DPO under this constructed preference dataset.
Table 3: Ablation on preference data: comparison between MAT-SFT, MAT-SFT-DPO, and SPORT on the GTA benchmark.
@@ 4.3.2 Applying DPO to MAT Preference Data
4.4 GAIA Results
In Table 4, we report the performance of Sport Agent on the GAIA validation set. Sport Agent achieves best results among agents that use open-source models, surpassing the best-performing open-source model, Qwen2-VL-7B, about $11\%$ on AnsAcc. The consistent improvements across different levels underscore the efficacy of our online self-exploration framework. Furthermore, Sport Agent demonstrates significant gains over the SFT-tuned controller, with an improvement about $4\%$ over MAT-Qwen2-VL-7B. However, when compared to agents leveraging closed-source models such as GPT-4, Sport Agent exhibits a slight performance gap. We attribute this discrepancy to the larger model sizes and more extensive training data available to closed-source models.
Table 4: Results on the validation set of the GAIA benchmark
Method Controller AnsAcc Level 1 Level 2 Level 3
Sibyl Agent GPT-4-turbo 29.70 43.40 27.90 7.70
Warm-up Act GPT-4-turbo 17.60 30.20 15.10 0.00
HF Agent GPT-4o 33.40 47.17 31.40 11.54
HF Agent GPT-4o mini 26.06 33.96 27.91 3.84
HF Agent LLaVA-NeXT-8B 3.64 9.43 1.16 0.00
HF Agent InternVL2-8B 4.85 7.55 4.65 0.00
HF Agent MiniCPM-V-8.5B 7.27 13.21 5.81 0.00
HF Agent Qwen2-VL-7B 9.70 16.98 8.14 0.00
T3-Agent MAT-MiniCPM-V-8.5B 15.15 26.42 11.63 3.84
T3-Agent MAT-Qwen2-VL-7B 16.97 26.42 15.12 3.84
Sport Agent Tuned Qwen2-VL-7B 20.61 35.85 16.28 3.84
@@ 4.4 GAIA Results
4.5 Statistic
We aggregated step-wise preference data from all the $4$ iterations, each with $d=500$ and a sampling size of 5, resulting in a total of 16K samples. We analyze the differences between the $chosen$ and $rejected$ step-wise data in terms of code error rate, tool selection, and content variations.
Tool Distribution. We analyze the distribution of tools in the chosen and rejected steps by examining their frequency of occurrence, as shown in Figure 3. The results highlight distinct differences in tool usage between the two groups.
@@ 4.5 Statistic
Figure 4: Example of constructed preference data.
[…]
Table 5: Comparison of code error rates between chosen and rejected steps.
chosen rejected
Code error rate 18.35% 81.94%
@@
Verifier BLEU-1 ($\downarrow$) BLEU-2 ($\downarrow$) BLEU-3 ($\downarrow$) BLEU-4 ($\downarrow$)
@@
Content Difference. We compared the BLEU scores of steps selected by our verifier and those selected randomly. BLEU scores measure the similarity between different sequences, with lower scores indicating higher discrimination. As shown in Table 6, our verifier consistently achieved lower BLEU scores across all n-grams compared to random selection. Specifically, for BLEU-1, BLEU-2, BLEU-3, and BLEU-4, our verifier’s scores were $0.30$, $0.18$, $0.14$, and $0.11$, respectively, while random selection yielded higher scores of $0.53$, $0.41$, $0.36$, and $0.34$. The results demonstrate that our verifier selects more distinct and differentiated steps, enhancing the quality of the chosen steps.
Figure 5: Comparisons between SFT Agent and our SPORT Agent.
Table 7: Average scores from humans.
@@
Reasonableness Naturalness Tool Content Code
@@ 4.6 Data Quality
The evaluation was performed for tasks and trajectories, based on five criteria: (1) ‘Reasonableness’ to evaluate whether the generated tasks are infeasible; (2) ‘Naturalness’ to evaluate whether the generated tasks are natural; (3) ‘Code’ to evaluate the accuracy of code in solutions; (4) ‘Tool’ to assess the appropriateness of tool selection; and (5) ‘Parameter’ to assess the correctness of parameter passing. Participants provided scores ranging from 1 to 10, with higher scores indicating better performance. As shown in Table 7, the average scores for our task and framework exceeded 8, demonstrating the validity of the preference data collected by our approach.
To further assess the effectiveness of the verifier in generating preference data, we conducted an additional study with a separate group of 20 researchers. Each participant was asked to evaluate 50 steps sampled in parallel by making preference selections. The preferences were determined using the same three criteria: code accuracy, tool selection appropriateness, and parameter passing correctness. We measured the agreements between the verifier’s preferences and those of the human participants, denoted as ‘Preference’ in Table 7. The results revealed an $82\%$ overlap. This high level of agreement validates the reliability of the verifier in capturing human-like preferences.

### v4

@@ 3.4 Preference Tuning
Objective. In one iteration, we may generate multiple tasks and construct preference data for them. After that, we denote the obtained preference data set as $\mathcal{D}=\{(x_{i},a_{i}^{pre},a_{i}^{dis})\}$. We choose the flexible DPO algorithm,
$\begin{aligned} \mathcal{L}(\theta)=-\mathbb{E}_{(x_{i},a_{i}^{pre},a_{i}^{dis}))\sim\mathcal{D}}[\log\sigma(\beta\log\frac{\pi_{\theta}(a_{i}^{pre}|x_{i})}{\pi_{ref}(a_{i}^{pre}|x_{i})}-\beta\log\frac{\pi_{\theta}(a_{i}^{dis}|x_{i})}{\pi_{ref}(a_{i}^{dis}|x_{i})})],\end{aligned}$ (3)
@@
Training Scheme. The proposed tool usage exploration is performed after an SFT stage for controllers, since the effectiveness of self-exploration requires the controller to have the ability to generate accurate actions. The SFT stage is the same as MAT gao2024multimodalagenttuningbuilding , where 20K trajectories are used to align the agent controller (Qwen2VL-7B in practice) with desirable outputs. In the self-exploration stage, we use preference tuning to update Qwen2-VL. The preference tuning process is summarized in  Algorithm˜1.
Algorithm 1 Training process in SPORT
0:  Seed of tasks, initial agent controller $\pi_{\theta}$ after SFT, and $\pi_{ref}=\pi_{\theta}$. Preference data $\mathcal{D}=\emptyset$.
0:  Updated agent controller $\pi_{\theta^{*}}$.
@@
6:     Sample $n$ actions $\{t_{i}^{1},c_{i}^{1},\ldots,t_{i}^{n},c_{i}^{n}\}$ based on the history $h_{i}$, and execute them to obtain results $\{o_{i}^{1},\ldots,o_{i}^{n}\}$.
7:     Select the best action $\{t_{i}^{\star},c_{i}^{\star}\}$.
8:     Construct n1n-1 preference pairs, and add them into $\mathcal{D}$.
@@ 4.1 Setting
The performance of the proposed SPORT approach is assessed on the GTA (wang2024gta, ) and GAIA (mialon2023gaia, ) benchmarks. Results are compared against agents powered by both closed-source models (e.g., GPT-4, GPT-4o, Claude3) and open-source models, including LLaMA-3-70B-instruct (dubey2024llama, ), Qwen1.5-72B-chat (bai2023qwen, ), LLaVA-NeXT-8B (liu2024llavanext, ), InternVL2-8B (chen2024far, ), Qwen2-VL-7B (wang2024qwen2, ), and MiniCPM-V-8.5B (yao2024minicpm, ). Specifically, we perform direct comparisons with leading agents, such as Lego Agent (AgentLego_Contributors_AgentLego_Open, ) and Warm-up Act Agent (mialon2023gaia, ). As a baseline, we use the Huggingface Agent (HF Agent) (huggingfaceagent, ), which operates with the same toolset as the SPORT Agent . We first evaluate these agents on two benchmarks, then assess the quality of the produced preference data, and finally show several visualization examples to demonstrate the effectiveness of our method.
[…]
@@
Metric. Following existing methods wang2024gta ; gao2024multimodalagenttuningbuilding , we assess agent performance using three key metrics: AnsAcc, ToolAcc, and CodeExec for the GTA benchmark. AnsAcc gauges the accuracy of predicted answers. ToolAcc evaluates the correctness of tool usage and the quality of answer summaries. CodeExec measures the percentage of generated code that executes without errors. In the GAIA benchmark, we focus on measuring AnsAcc at its three levels.
Table 1: Results on two benchmarks: GTA and GAIA. The bold results represent the best performance compared to the open-source models.
Method Controller GTA GAIA
ToolAcc CodeExec AnsAcc Level 1 Level 2 Level 3 AnsAcc
Closed-source Controller
Lego Agent GPT-4 - - 46.59 - - - -
Lego Agent GPT-4o - - 41.52 - - - -
Warm-up Agent GPT-4-turbo - - - 30.20 15.10 0.00 17.60
HF Agent GPT-4o 63.41 95.12 57.05 47.17 31.40 11.54 33.40
HF Agent GPT-4o-mini 56.10 100.00 57.69 33.96 27.91 3.84 26.06
Open-Source Controller
HF Agent LLaVA-NeXT-8B 14.97 25.08 14.10 9.43 1.16 0.00 3.64
HF Agent InternVL2-8B 36.75 52.18 32.05 7.55 4.65 0.00 4.85
HF Agent MiniCPM-V-8.5B 36.59 56.10 33.97 13.21 5.81 0.00 7.27
HF Agent Qwen2-VL-7B 44.85 65.19 42.31 16.98 8.14 0.00 9.70
T3-Agent MAT-MiniCPM-V-8.5B 65.85 80.49 52.56 26.42 11.63 3.84 15.15
T3-Agent MAT-Qwen2-VL-7B 64.63 84.32 53.85 26.42 15.12 3.84 16.97
Ours
SPORT Agent Tuned-Qwen2-VL-7B 72.41 91.87 60.26 35.85 16.28 3.84 20.61
@@ 4.2 GTA Results
The results on the GTA benchmark are shown in Table˜1, where key metrics including AnsAcc, ToolAcc, and CodeExec are reported. Our agent surpasses the Lego agent that utilizes closed-source models (e.g., GPT-4 and GPT-4o), as well as the HF agent that uses closed-source models and open-source models (e.g., InternVL2-8B), showcasing the ability of our SPORT Agent to tackle complex tasks with greater efficiency. A comparison between agents through SFT (i.e., T3-Agent) and our SPORT Agent demonstrates the effectiveness of our online self-exploration framework and the advantages of our Step-wise optimization approach. Our method has about $7\%$ improvements on the final accuracy, since it calls more suitable tools ($8\%$ improvements) and reduces code error ($7\%$ improvements). Compared with the HF agent using GPT-4o and GPT-4o mini, our agent achieves higher ToolAcc and comparable CodeExec. This indicates that the proposed SPORT method improves the planning and reasoning capabilities of agents again.
4.3 GAIA Results
In Table˜1, we report the performance of SPORT Agent on the GAIA validation set. SPORT Agent achieves best results among agents that use open-source models, surpassing the best-performing open-source model, Qwen2-VL-7B, by about $11\%$ on AnsAcc. The consistent improvements across different levels underscore the efficacy of our online self-exploration framework. Furthermore, SPORT Agent demonstrates significant gains over the SFT-tuned controller, with an improvement of about $4\%$ over MAT-Qwen2-VL-7B. However, when compared to agents leveraging closed-source models such as GPT-4, SPORT Agent exhibits a slight performance gap. We attribute this discrepancy to the larger model sizes and more extensive training data available to closed-source models.
Table 2: Ablation on iteration step $d$ in the GTA benchmark.
@@ 4.3 GAIA Results
Table 3: Ablation on preference data: MAT-SFT vs. MAT-SFT-DPO vs. SPORT on the GTA benchmark.
@@
4.4 Ablation
Effectiveness of Iteration Step Size
We conduct an ablation study on the GTA benchmark to investigate the impact of the iteration step size $d$ that denotes the number of used trajectories in each iteration. We set $d\in\{200,500,1000\}$, adjusting the number of iterations to (5, 2, 1), respectively, to ensure a total of 1000 trajectories are processed in each setting. As shown in Table 2, setting $d=500$ yields the best overall performance across all metrics. When $d=1000$, the model sees all tasks in a single pass, which limits adaptability and leads to a slight drop in answer accuracy and execution success. In contrast, using a smaller step size ($d=200$) results in less diverse updates per iteration, which may reduce robustness. These results suggest that a moderate value of $d$ offers a good balance between update frequency and data diversity, leading to more stable and effective training.
Comparison with Static Preference Data
We conduct ablation experiments to compare our method (online exploration) with DPO on static tasks. In doing so, we directly apply DPO to the MM-Traj dataset gao2024multimodalagenttuningbuilding . Specifically, we treat the MM-Traj data (GPT-4o generated) as “preferred” samples and synthetically generate an equal number of “dispreferred” examples via the MAT-SFT model, matching the total volume of our SPORT preference data. We then fine-tune MAT-SFT using DPO under this constructed preference dataset.
Training with Different Base Models
To evaluate the generalizability of SPORT, we apply it to different base models. We compare four configurations on the GTA benchmark: Qwen2-VL-7B (base model), MAT-Qwen2-VL-7B (base model with MAT applied), SPORT-Qwen2-VL-7B (SPORT applied directly to base model), and SPORT-MAT-Qwen2-VL-7B (SPORT applied to MAT-tuned model).
Table 4 presents the answer accuracies. Applying SPORT directly to Qwen2-VL-7B improves accuracy from 42.31% to 55.13% (+12.82%), demonstrating SPORT’s effectiveness as a standalone method. When applied to MAT-tuned baseline, SPORT achieves the highest accuracy of 60.26% (+5.13% over MAT alone), indicating that SPORT can effectively complement existing tuning methods. These results confirm that SPORT is a flexible approach that enhances agent performance both independently and in combination with other preference optimization methods.
Sensitivity to Task Quality and Diversity
To assess SPORT’s robustness to the quality and diversity of synthetic tasks in early-stage self-exploration, we conduct an ablation study where the in-context examples are reduced from 20 to 5 and the task seed pool is narrowed from 425 to 100. This setup mimics a low-diversity scenario during early-stage exploration.
As shown in Table 5, this reduction leads to a moderate drop in performance (60.26 $\rightarrow$ 58.33). However, the result still significantly outperforms the SFT baseline (58.33 vs. 53.85), demonstrating that SPORT maintains strong performance even under constrained task diversity. These findings indicate that while task quality and diversity do impact SPORT’s effectiveness, the method exhibits robustness to variations in the synthetic task generation process.
Table 4: Answer accuracies (%) on GTA benchmark with different base models.
Model AnsAcc (%)
Qwen2-VL-7B 42.31
MAT-Qwen2-VL-7B 53.85
SPORT-Qwen2-VL-7B 55.13
SPORT-MAT-Qwen2-VL-7B 60.26
Table 5: Impact of task diversity on the GTA benchmark.
Method AnsAcc (%)
MAT-Qwen2-VL-7B 53.85
SPORT w/ 5 from 100 seeds 58.33
SPORT w/ 20 from 425 seeds 60.26
4.5 Statistic
We aggregate step-wise preference data from all iterations, each with $d=500$ and a sampling size of 5, resulting in a total of 16K samples. We analyze the differences between the chosenchosen and rejectedrejected step-wise data in terms of code error rate, tool selection, and content variations.
@@ 4.5 Statistic
Tool Distribution
We analyze the distribution of tools in the chosen and rejected steps by examining their frequency of occurrence, as shown in Figure˜3. The results highlight differences in tool usage between the two groups.
@@
Code Error Rate
We compared the code execution status of chosen and rejected steps by measuring the proportion of execution results (Observations) that contained code errors. The rejected steps exhibit a significantly higher error rate (81.94%) compared to the chosen steps (18.35%). This indicates that our step-wise preference data favors code that executes successfully. Consequently, this preference also leads to the improvement in the code accuracy of the SPORT Agent, as shown in Table˜1 ($CodeExec\ \ 84.32\%\rightarrow 91.87\%$).
Content Difference
We compared the BLEU scores of steps selected by our verifier and those selected randomly. BLEU scores measure the similarity between different sequences, with lower scores indicating higher discrimination. As shown in Table˜7, our verifier consistently achieved lower BLEU scores across all n-grams compared to random selection. Specifically, for BLEU-1, BLEU-2, BLEU-3, and BLEU-4, our verifier’s scores were $0.30$, $0.18$, $0.14$, and $0.11$, respectively, while random selection yielded higher scores of $0.53$, $0.41$, $0.36$, and $0.34$. The results demonstrate that our verifier selects more distinct steps, enhancing the quality of the chosen steps.
@@
Verifier B1 ($\downarrow$) B2 ($\downarrow$) B3 ($\downarrow$) B4 ($\downarrow$)
@@
Table 7: Average scores from humans on data quality.
@@
Reasonable Natural Tool Content Code
@@ 4.6 Data Quality
The evaluation was performed for tasks and trajectories, based on five criteria: (1) ‘Reasonableness’ to evaluate whether the generated tasks are infeasible; (2) ‘Naturalness’ to evaluate whether the generated tasks are natural; (3) ‘Code’ to evaluate the accuracy of code in action; (4) ‘Tool’ to assess the appropriateness of tool selection; and (5) ‘Parameter’ to assess the correctness of parameter passing. Participants provided scores ranging from 1 to 10, with higher scores indicating better performance. As shown in Table˜7, the average scores for our task and framework exceeded 8, demonstrating the validity of the preference data collected by our approach.
To further assess the effectiveness of the verifier in generating preference data, we conducted an additional study with a separate group of 20 researchers. Each participant was asked to evaluate 50 steps sampled in parallel by making preference selections. The preferences were determined using the same three criteria: code accuracy, tool selection appropriateness, and parameter passing correctness. We measured the agreements between the verifier’s preferences and those of the human participants, denoted as ‘Preference’ in Table˜7. The results revealed an $82\%$ overlap. This high level of agreement validates the reliability of the verifier in capturing human-like preferences.
Figure 4: Example of constructed preference data.
Figure 5: Comparisons between SFT Agent and our SPORT Agent.